OpenStudy (anonymous):

Which expression is equivalent to sec^2xcot^2x? A)sin^2x B)csc^2x C)1\cos^2x D)1\sin^2x

OpenStudy (ipwnbunnies):

\[\sec x = \frac{1}{\cos x}\] \[\cot x = \frac{\cos x}{\sin x}\] These are the two identities you'll need

OpenStudy (ipwnbunnies):

\[\sec^2 x \cot^2 x\] = \[(\frac{1}{\cos x})^2 (\frac{\cos x}{\sin x})^2\]